OpenStudy (anonymous):

A cylinder and a rectangular prism have the same volume and same height. The base of the prism is a square with side length 8 cm. What is the approximate radius of the cylinder? A. 2.5 cm B. 4.0 cm C. 4.5 cm D. 20.4 cm

OpenStudy (anonymous):

help

OpenStudy (anonymous):

pls ssss fast helpp

OpenStudy (anonymous):

64 = r^2*pi

OpenStudy (anonymous):

i am out of breath

OpenStudy (anonymous):

64/pi = r^2

OpenStudy (anonymous):

8/sqrt(pi) = r 4.51=r

OpenStudy (anonymous):

k

OpenStudy (anonymous):

C are you running or something?

OpenStudy (anonymous):

no i am in a marothon

OpenStudy (anonymous):

hahaha